OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5?

OpenStudy (anonymous):

are either of these correct y = -5x + 22 y = 5x + 22

OpenStudy (hybrik):

As I done before, you plug in the points to \[Y=-1/5x + B\]

OpenStudy (anonymous):

wait what are you supposed to plug in for B

OpenStudy (hybrik):

which would be \[7=-1/5*-3 + B\]

OpenStudy (hybrik):

sher, your trying to find B, the y intercept

OpenStudy (hybrik):

Your solving for B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you solve it

OpenStudy (hybrik):

As I put, plug in the X value for X and Y value for y in formula as shown above

OpenStudy (anonymous):

so I've got b=1/5*3+7 what's next?

OpenStudy (hybrik):

Simplify that

OpenStudy (anonymous):

b=7.6

OpenStudy (anonymous):

so is that supposed to be the correct answer

OpenStudy (hybrik):

No, getting the y intercept, plug in "B" to \[Y = -1/5x + B\]

OpenStudy (anonymous):

is that one of these

OpenStudy (hybrik):

No, is there more choices

OpenStudy (anonymous):

y = -5x + 22 y = 5x + 22

OpenStudy (anonymous):

help me plz